OpenStudy (mendicant_bias):

ODE question asking for a general solution and an interval on which the general solution is defined. Posted as soon as the OS thread interface starts working again.

OpenStudy (mendicant_bias):

Alright, http://i.imgur.com/Rp7Sdnc.png

OpenStudy (mendicant_bias):

That's the prompt (Not going to write it out in LaTeX anymore, lol, time is of the essence)

OpenStudy (anonymous):

Okay, seems pretty straightforward. You know the kind of equation and how to approach it?

OpenStudy (mendicant_bias):

Thinking about it, half of the time I'm doing this, particularly because most of this stuff I've learned in the last four hours are so, my memory of application is horrific. Give me a second to try to remember it on my own.

OpenStudy (anonymous):

kk, no problem :)

OpenStudy (mendicant_bias):

Alright, so this is a nonhomogeneous, linear ODE with non-constant coefficients. First order linear ODE's when put into proper form are best solved with either separation of variables or using an integrating factor, IIRC.

OpenStudy (anonymous):

Yep, ya figure its one of the two.

OpenStudy (mendicant_bias):

\[x \frac{dy}{dx}+(4x+1)y=x^2e^{4x}\]Wait, this isn't a Bernoulli DE, is it? I don't remember the general form of one but for some reason this is reminding me of a Bernoulli DE.

OpenStudy (mendicant_bias):

Nevermind, nevermind

OpenStudy (anonymous):

Bernoulli would be a higher power of y on the other side of the equation.

OpenStudy (mendicant_bias):

Yeah. Alright, then, either separation of variables or using an integrating factor, one minute.

OpenStudy (mendicant_bias):

Going to first put it in the proper form to use an integrating factor by getting y prime to have a coefficient of one.

ganeshie8 (ganeshie8):

f(x) y' + g(x) y = h(x) linear equation

OpenStudy (mendicant_bias):

\[\frac{dy}{dx}+\frac{4x+1}{x}y=xe^{4x}\]

OpenStudy (anonymous):

Mhm. So definitely the form where you'd have an integrating factor.

OpenStudy (mendicant_bias):

Alright, now have to assume that the LHS is the derivative of the product of some mystery function mu, and y.

OpenStudy (mendicant_bias):

Oh, nevermind, I need to multiply the whole function by mu first in order to do that.

OpenStudy (mendicant_bias):

\[\mu \frac{dy}{dx}+\mu \frac{4x+1}{x}y=\mu xe^{4x}\]

OpenStudy (mendicant_bias):

\[\mu '=\mu P(x) = \mu \frac{4x+1}{x}\]

OpenStudy (mendicant_bias):

\[\frac{\mu '}{\mu}=\frac{4x+1}{x}\]

OpenStudy (mendicant_bias):

\[\int\limits_{}^{}\frac{\mu '}{\mu} d \mu=\int\limits_{}^{}\frac{4x+1}{x}dx\]

OpenStudy (mendicant_bias):

\[\ln(\mu)+c_{1}=4x+\ln(x)+c_{2}\]

OpenStudy (mendicant_bias):

I know I need to isolate mu by exponentiating both sides, but I start getting confused about how to properly handle these constants of integration, if at all; I know sometimes they are amalgamated together into a single constant.

OpenStudy (unklerhaukus):

\[\ln(\mu)+c_{1}=4x+\ln(x)+c_{2}\\ \ln(\mu)=4x+\ln(x)+c_{3}\\ \mu=e^{4x+\ln x+c_3}\\ \quad=Ce^{4x+\ln x}\]

ganeshie8 (ganeshie8):

you just need some \(\mu\) so set the constant to 0

OpenStudy (mendicant_bias):

Alright, yeah, now I remember how that works, thank you. One sec.

ganeshie8 (ganeshie8):

I dont think your professor really wants you to carry out this process in every single problem. you could simply use \[IF = e^{\int p~dx}\]

OpenStudy (mendicant_bias):

He does.

OpenStudy (unklerhaukus):

\[\frac{dy}{dx}+p(x)y=q(x)\]\[\mu(x) =e^{\int p(x)\,dx}\] \[\frac{d}{dx}(\mu(x) y)=\mu(x) q(x)\]

OpenStudy (mendicant_bias):

We're not allowed to memorize that formula, and we have to work it that way every single time.

OpenStudy (mendicant_bias):

Which is ridiculous and sucks, but it's the way he does it. The book also implores us repeatedly in more than one section, after it's given a definitive proof for how to get there, to do it from the ground up. Every single time.

OpenStudy (anonymous):

Because that's soooo much more of a complicated way to think about it. Once you get the coefficient of y' to be 1 in an equation, all you need to do is take the integral of the function of x multiplying y and raise that as an exponent of e. Given an equation in the form: y' + P(x)y = Q(x), the integrating factor is equal to \[\mu = e^{\int\limits_{}^{}P(x)dx }\] For you, P(x) = 4 + 1/x, so you get \[\mu = e^{4x + lnx} = xe^{4x}\] The only constant of integration you need is the last one. Now is when you worry about multiplying by mu. In the end, your integral will always end up being \[y \mu = \int\limits_{}^{}\mu Q(x)dx\] Your Q(x) is \(xe^{4x}\), so \[yxe^{4x} = \int\limits_{}^{}x^{2}e^{8x}dx\] Wait, you have to do it the long way every time? Lame :(

ganeshie8 (ganeshie8):

thats a good thing to enforce on grade school kids but im sure that must be frustrating for grown up students lol

OpenStudy (mendicant_bias):

Yeah, it is, heh. Because these are freaking timed exams.

OpenStudy (unklerhaukus):

The final solution to a first order DE, should have one constant in it. This means you dont have to worry about constants in the integrating factor, because you'll get a constant of integration later anyway

OpenStudy (mendicant_bias):

Alright, let me try to pick up where we were, got lost in the convo.

OpenStudy (mendicant_bias):

\[\mu = Ce^{\ln(x)+4x}\]

ganeshie8 (ganeshie8):

Right, give C your favorite value

OpenStudy (unklerhaukus):

\[\int\limits_{}^{}\frac{\mu '}{\mu} d x=\int\limits_{}^{}\frac{4x+1}{x}dx\\ \ln(\mu)=4x+\ln(x)\\ \mu = e^{4x+\ln(x)}\]

OpenStudy (mendicant_bias):

Lul "your favorite value"

OpenStudy (mendicant_bias):

Yeah, now just remembering how to finish the problem out

OpenStudy (anonymous):

Favorite value, hehe \[\mu = (banana)e^{\ln(x) + 4x}\]

ganeshie8 (ganeshie8):

finding the domain might be tricky

OpenStudy (mendicant_bias):

Alright, so plugging into the general form of the equation modified to have mu:

ganeshie8 (ganeshie8):

yes if banana is not a function of x :P

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Banana is an unknown constant D:

OpenStudy (mendicant_bias):

\[e^{\ln(x)+4x}\frac{dy}{dx}+e^{\ln(x)+4x}\frac{4x+1}{x}y=e^{\ln(x)+4x}xe^{4x}\]

OpenStudy (mendicant_bias):

(Jesus what is this I can't even)

OpenStudy (mendicant_bias):

This is such a gross looking problem :C https://33.media.tumblr.com/f8bdf5093c849397f4aca47fbfcf7b9c/tumblr_ndlex59ktC1rdfsmyo1_500.gif

OpenStudy (anonymous):

Yeah, I'll always add cheesy humor to my math. If I need imaginary solutions, I consider chair an imaginary solution. And thats the mess you get when you have to do it all th way through, step-by-step.

OpenStudy (mendicant_bias):

WELL

OpenStudy (mendicant_bias):

I think what I'm going to do if we get something like this on the exam is do like 90% of the workings up until this point....and then just sort of try to jump to the final result, near the end. lol.

OpenStudy (unklerhaukus):

Use the reverse product rule\[\mu(x)\frac{dy}{dx}+\mu(x)p(x)y=\mu(x)q(x)\] \[\frac{d}{dx}\Big(\mu(x) y\Big)=\mu(x) q(x)\]

OpenStudy (mendicant_bias):

Alright, so, I got that, I just again am bad at remembering how to put these solutions in their proper forms once I have them. So what we have is \[e^{\ln(x)+4x}y=e^{\ln(x)+4x}xe^{4x}\]

OpenStudy (mendicant_bias):

(Double checking that)

ganeshie8 (ganeshie8):

*derivative* symbol is missing on left side

OpenStudy (mendicant_bias):

Nah, I dun goofed, I better be careful, and thanking you for mentioning that, I see what I need to do, I just didn't realize it for a moment, integrate both sides.

OpenStudy (mendicant_bias):

\[\int\limits_{}^{}\bigg[\frac{d}{dx}\bigg(e^{\ln(x)+4x}y\bigg)dx\bigg]=\int\limits_{}^{}e^{\ln(x)+4x}xe^{4x}\]

OpenStudy (mendicant_bias):

\[\int\limits_{}^{}e^{\ln(x)+4x}xe^{4x}=\int\limits_{}^{}e^{\ln(x)}e^{4x}xe^{4x}=\int\limits_{}^{}x^2e^{8x}\]

OpenStudy (mendicant_bias):

Integration by parts? Really? This problem would take *forever* on a test, lord.

ganeshie8 (ganeshie8):

it should take less than 10 seconds to come this far if you stop worrying about ur professor

OpenStudy (mendicant_bias):

Lul

OpenStudy (anonymous):

Yeah, this whole process is fast really. And even the by parts is fast. You know the tabular method for by parts, correct?

OpenStudy (mendicant_bias):

Yeah, well, at least I know the technique, I think I'll drop it here for the moment, if I conceptually get how to do it, I can leave the algebra for the test. For now, I'm running out of time on the more crucial thing, the concepts. T-minus 19 hours, 30 minutes. No, I mean, I've heard of it, but I'm not really familiar with it and don't use it. Let me put it this way: I posted this problem 40 minutes ago and *still* haven't solved it, granted I wasn't sure/didn't entirely remember how to 40 minutes ago, lol.

OpenStudy (mendicant_bias):

>puts hands under butt, waits for tabular method fancy demonstration

OpenStudy (anonymous):

Lol, you sure it wont be cramming extra info? :P

OpenStudy (mendicant_bias):

Yeah, you're right, I'll leave it as is, new question, new concept coming up.

OpenStudy (anonymous):

Sounds good.

OpenStudy (mendicant_bias):

(Also, interval on which solution is defined, I'm going to find an easier/quicker one to do that with.) HEY, EVERYBODY LOOKING AT THIS PROBLEM:

OpenStudy (mendicant_bias):

Do you think this was the right approach, or should I have attempted to solve this DE in another manner/using a different technique?

OpenStudy (anonymous):

Yeah, method of integrating factors. Just would be nice if you didnt have to go through all the in between mess.

OpenStudy (mendicant_bias):

Alright, cool. Hi-ho, silver, next problem. Thanks, folks.

ganeshie8 (ganeshie8):

can you post the interval in which the solutions are defined if psble

OpenStudy (mendicant_bias):

^I'm not sure, one minute, focusing on the other problem